# Words as Trigger Points in Social Media Discussions: A Large-Scale Case Study about UK Politics on Reddit

### v3

@@ 3. Temporal
We estimate the trigger words’ treatment effect in 60 separate experiments—12 experiments for H1 (Table 3) and 48 experiments for H2 (Table 5). For each experiment, we prepare the data in the same way. The treatment is the first mention of a trigger word in a discussion thread, i.e., there is only one well-defined treatment per thread. To test H1, we count the number of comments for each discussion thread 30 minutes before and 30 minutes after the treatment in the treatment group and the placebo treatment in the control group. We then calculate the share of comments before and after the intervention. For example, if there are 8 comments during the 30 minutes before the trigger word and 12 comments during the 30 minutes after the trigger word, we normalize the data of this discussion thread to 40% before intervention and 60% after intervention. For experiments with H2, we count the number of comments expressing animosity in a thread and then normalize the data as a share of animosity per thread before and after the trigger word.
We estimate the causal effect in an OLS-regression framework. More specifically, we specify the DiD with the two-way fixed effects estimator (Angrist and Pischke 2009):
[…]
where proportion is the per-thread-normalised target feature and the interaction term is given by
[…]
@@
Since the DiD estimator subtracts the difference before and after intervention in the control from the equivalent in the treatment group, no further control variables are necessary—if the parallel trends assumption holds (Roth et al. 2023). It requires that in the absence of treatment, the difference between the outcomes of the treated and control groups would follow the same trajectory over time. We investigate the plausibility of the assumption in section 4.3.
3.4 Features
We rely on several features to examine our initial hypotheses and study the impact of trigger words in online discussions. We use comment counts for the first hypothesis (H1) about user engagement. The second hypothesis (H2) requires a measure of animosity. To identify relevant features that can operationalize animosity, we build on various language models that follow the RoBERTa architecture (Liu et al. 2019) and are trained in social media corpora, specifically X (Twitter). Even though comments on Reddit tend to be longer than on X, the posts on both platforms share many similarities, e.g., the use of slang, emojis, or unstructured text. Hence, we expect the trained models to perform well in either context (Priya et al. 2019).
User engagement (Hypothesis 1)
@@ 3.4 Features
Our first hypothesis suggests that trigger words lead to higher user engagement. We rely on the number of comments to investigate this and expect an uptick in message frequency after the trigger words are mentioned for the first time in a thread.
Animosity (Hypothesis 2)
@@
Hypothesis 2 suggests that trigger words cause more animosity in a debate. We operationalize this with controversial comments, negative sentiment, anger, and different forms of hate speech.
Controversiality
@@
We examine whether a comment is controversial using the Reddit controversiality feature. Reddit offers a measure that identifies a comment as controversial if it receives similar amounts of upvotes and downvotes. This metric is based on the users’ approval, so care must be taken when considering already controversial subreddits. We expect trigger words to cause more controversiality.
@@
Figure 2: Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
Sentiment analysis
@@
For sentiment analysis, we use twitter-roberta-base-sentiment-latest (Loureiro et al. 2022) which identifies a comment as negative, neutral, or positive. This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal, Farra, and Nakov 2019). As we expect trigger words to cause more negative sentiments, we use sentiment analysis to investigate this, similarly to work using it to gain insights into how controversial topics such as vaccination (Melton et al. 2021), politics (Guimaraes et al. 2019), and gender biases (Marjanovic, Stańczak, and Augenstein 2022) are discussed online.
Negative emotions
@@
We leverage twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al. 2022) to assign one or more emotions to each tweet. This model is trained on the Semeval 2018 Affect in Tweets task (Mohammad et al. 2018), covering 11 different emotions. [^6] Detailed list of emotions can be found in Appendix A.1) For our analysis, we focus on anger specifically, and we expect to observe more anger after a trigger word.
Figure 3: Distribution of negatively charged comments for the treatment and control (space) groups of Vaccine as a trigger word.
Hate speech
@@
Finally, we use the twitter-roberta-base-hate-multi-class model (Antypas and Camacho-Collados 2023) to detect hate speech. It is trained on a combination of 13 different hate speech Twitter datasets and can predict hate towards seven target groups. [^7] Detailed list of categories can be found in Appendix A.2 ) The inclusion of hate speech detection as a feature is motivated by previous research revealing hateful trends in topics such as politics (Rieger et al. 2021), women’s rights (Farrell et al. 2019), and health issues (Walter et al. 2022). Considering the connotations of some of the triggers, we specifically focus on sexism for Feminism/Feminist and racism for Rwanda. For NHS, Brexit and Vaccine we consider all hateful comments jointly.
3.5 Validation of Annotation Model
The primary aim of our research is to use simple and computationally efficient models that reliably extract signals relevant to our hypotheses. To confirm the validity and robustness of our findings, we implemented an additional validation procedure to ensure consistency across the classifiers employed in our analysis.
We sampled 100 Reddit comments from our data collection, with 20 comments selected from the subreddits associated with each trigger point (Table 2). The collected comments were independently annotated by three experienced NLP and CSS researchers for sentiment, emotion, and the presence of hate speech. [^8] We include the annotation guidelines in Appendix A.4. Each comment was assigned a label based on the majority vote, while disagreements were resolved through discussion. When considering the difficulty of the annotation task, the annotations present above-average agreement scores for sentiment and hate speech, with Fleiss’ $\kappa$ scores of 0.52 and 0.56, respectively. For the multi-label emotion annotation task, where annotators had to choose from a set of 12 labels,the Krippendorff’s alpha score (Krippendorff 2011) reached 0.24, which is in line with other similar multi-label tasks with a high number of labels (Mohammad et al. 2018; Antypas et al. 2022; Ousidhoum et al. 2019).
We fine-tune the base version of BERT-cased, XLNet, and RoBERTa for sentiment analysis, emotion recognition, and hate speech detection on the same datasets as the models we used for our experiments, as explained in Section 3.4 (Rosenthal, Farra, and Nakov 2017; Mohammad et al. 2018; Antypas and Camacho-Collados 2023) to ensure fairness. We further compare our results to those of five pre-trained models publicly available on HuggingFace. Specifically, two sentiment classifiers trained on Reddit data999 reddit-dBERT: https://huggingface.co/akshataupadhye/finetuning-sentiment-model-reddit-data, reddit-XLNET: https://huggingface.co/minh21/XLNet-Reddit-Sentiment-Analysis , hateBERT (Caselli et al. 2021) and one emotion classifier [^10] BERT-emotion: https://huggingface.co/ayoubkirouane/BERT-Emotions-Classifier. All models are evaluated in the aforementioned manually labelled dataset.
Table 4 presents the average micro F1-scores achieved by the models across each task. We opted for the micro F1-score as it provides a more realistic representation of the results, given the limited sample and heavily imbalanced nature of our dataset. [^11] Detailed dataset distribution is available in Appendix A.3 The Twitter-RoBERTa models achieved the highest scores on all tasks.These scores are matched by XLNet-base for sentiment analysis and XLNet for emotion detection, and the fine-tuned models RoBERTa and XLNet for hate speech detection. Overall, the tested models performed comparably well, except for two off-the-shelf sentiment classifiers that achieved scores below 50%.
Sentiment Emotion Hate
t-RoBERTa 0.76 0.50 0.88
RoBERTa 0.75 0.47 0.88
XLNet 0.76 0.45 0.88
BERT 0.74 0.40 0.85
reddit-dBERT 0.45 - -
reddit-XLNet 0.31 - -
BERT-emotion - 0.45 -
hateBERT - - 0.80
Random 0.30 0.20 0.07
Table 4: Average micro-F1 scores of the models tested for the sentiment analysis, emotion, and hate speech detection tasks. Random indicates a naive baseline that randomly selects labels.
4 Results
4.1 User engagement (Hypothesis 1)
@@
According to our first hypothesis, trigger points cause more comments. The DiD estimate reveals significant differences between the frequency of comments in the treatment and control groups for all our trigger words (Table 3).
@@
In the treatment-only subset, the difference between user engagement before and after the trigger word is apparent for all terms. It is the largest with Rwanda where we observe a 10%-points (Table 3) increase. To identify the causal effect of the trigger word, we compare the difference in the treatment group to the difference in the control group. In every setting, we observe a significant difference in the number of comments in the threads between the treatment and control subsets, with the frequency of comments in the treatment group increasing by a larger factor according to the DiD estimate. The lowest value is a 3.2%-point increase for the contextual comparison in Brexit. The largest increase is for NHS in the contextual control group, where we observe a 13.4% point increase.
@@
When examining comment frequency across various control groups, the impact of specific trigger words becomes more apparent. For instance, Figure 2 illustrates the distribution of comments related to the trigger Rwanda in the space-related experiment. It shows the overall number of messages across all threads. This trigger word leads to varying levels of engagement across different Reddit communities, with a notable increase in comments in identified communities where Rwanda is more controversial. Similar patterns are also observed in the contextual and temporal control groups for Rwanda (Appendix B, Figure 10).
Overall, our analysis of the influence of trigger words across different control groups indicates a clear increase in engagement for each trigger word analysed, confirming our first hypothesis.
@@
Table 5: Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
4.2 Animosity (Hypothesis 2)
@@
Our second hypothesis suggests that trigger words may increase the frequency of polarised messages and lead to more animosity in the discussion. [^12] An example of such a discussion is presented in Appendix E.
@@
Figure 4: Average controversiality per thread for the treatment and control (space) groups of Rwanda.
@@
Controversiality. We analyse the average controversiality of each thread and compare the proportion of controversial comments before and after the introduction of the trigger words. Our results (Table 5) indicate that, in general, there is an increase in controversial comments following each trigger word. Most terms display a statistically significant increase of DiD estimates when considering controversial messages in the treatment and control groups, with the biggest difference observed on the contextual comparison for Rwanda—a causal effect of 6.6% points.
However, significant differences in controversiality do not appear across all settings. Notably, there seems to be no increase in controversial comments for Feminism, and in the temporal and contextual settings for Vaccine and NHS, respectively.
@@
Figure 4 shows the percentages of controversial threads related to the trigger word Rwanda for the space experiment, before and after its occurrence. The data indicates a noticeable increase in thread controversiality caused by the trigger word, reflected by a DiD estimate of 2.3% points. The contextual and temporal control settings follow a similar trend, as documented in Appendix B, Figure 9.
(a) Rwanda
(b) Feminism
Figure 5: Distribution of hateful comments in the treatment and control groups (space) of the trigger points Rwanda and Feminism.
@@
Sentiment & emotion. We further investigate the effect trigger words have on sentiment and emotion. Based on our hypothesis, trigger words cause users to express more negatively charged emotions. In our analysis, we focus on the presence of negative sentiment and, more narrowly, the presence of anger in emotions.
@@
Table 5 indicates a clear trend for the negatively charged comments of the treatment and control groups for each trigger word. We observe a significant increase in comments conveying a negative sentiment after the trigger word appears in 11 of the 12 settings tested, with the sole exception being the contextual comparison for NHS. The largest effect in sentiment, with a DiD estimate of 5.5% points, appears for the Rwanda trigger word in the space and contextual experiments.
Let us consider the vaccine trigger, Figure 3, as an example. Here, we again plot the overall number of messages we observe. Both treatment and space control groups exhibit a similar pattern of negative sentiment density as time approaches the trigger word, with an increase in negative sentiment just before and after the trigger word is mentioned. However, the treatment group displays a noticeably higher volume of messages around the trigger term, suggesting that triggers amplify the expression of negative sentiment compared to the control condition.
@@
Similarly to the increase in negative sentiment, trigger words systematically cause more anger. We observe a positive causal effect in all conditions, with only 3 non-significant effects. Substantively, we measure a causal effect that ranges from 1.4% points more anger (Feminist) to 5.5% points (Vaccine contextual comparison). As another example, the overall number of comments expressing anger in the Vaccine treatment and space control groups increase significantly following the appearance of the trigger word, as shown in Figure 12, Appendix B. This is further confirmed by the DiD estimates, which indicate a significant difference in anger-related comments in 9 out of 12 cases.
@@
Hate speech. As a final indicator of animosity, we also consider hate speech and abusive language, more specifically any kind of toxicity for Brexit, NHS, and Vaccine, and the more domain-specific racism-related hate speech for Rwanda and sexism-related hate speech for Feminism. As seen from Table 5, the impact of trigger words regarding the occurrence of hateful or offensive comments is not as prevalent as anger and negative sentiment. We observe significant effects in those areas where we can count on a domain-specific hate speech measure, i.e., in the cases of Feminism (space and contextual comparison) and Rwanda (space group).
We illustrate our findings in a visualization, where we consider all messages irrespective of the thread in which they are posted. We observe a distinct pattern, particularly in the context of Rwanda and Feminism, where the volume of hateful content increases after the trigger. Figure 5 reveals the prevalence of racist and sexist comments within the treatment groups when compared to their control group in the space experiment.
4.3 Robustness of Results
Testing the Parallel Trends Assumption.
The DiD estimator relies on the assumption of parallel trends: without the treatment, the treated and control groups would follow the same trajectory over time. While there are no formal guarantees for the assumption to hold, there are means to argue in its favour. The first argument is a theoretical one. In neither of our experimental designs did we have reason to believe that the frequency of expressing animosity would not follow the same trend without the trigger word. Whether comparing treatment and control groups based on different Reddit fora (spaces), different words (contextual), or points in time (temporal), we suggest that without the respective trigger word, the relative difference between treatment and control group would remain the same.
We implement event studies for each experiment to further empirically support our argument. The central idea is to model the respective outcome variable as it unfolds over time, which allows us to assess the dynamics of treatment impact before and after the intervention. Plotting the treatment effects at different time points before the treatment, the event study helps visualize whether the pre-treatment trends in the treatment and control groups are indeed parallel. If the trends are not parallel before the intervention, the parallel trends assumption is violated, indicating potential bias in the DID estimator (Roth et al. 2023).
[…]
Figure 7 charts the time-wise means over all 60 event studies. If the parallel trends assumption holds, the lines before the treatment at time = 0 should all remain on the horizontal line where the treatment effect = 0. Indeed, this is largely what we find. The mean treatment effect is close to 0 before the trigger word is mentioned for the first time. After the intervention, it is clearly above the horizontal 0-line. [^13] See Appendix for figures that include individual event studies for all experiments from H1 and H2.
Placebo Treatment.
To further validate our Difference-in-Differences (DiD) estimates, we implement a placebo treatment analysis. The core logic of this approach is to check whether the estimated treatment effects could be driven by other trends earlier or later than the actual intervention.
In the placebo analysis, we systematically shift the assumed treatment time in 10-minute intervals, ranging from 30 minutes before to 30 minutes after the actual intervention, i.e., the comment with the first mention of the trigger word in a discussion thread. At each placebo point, we re-estimate all DiD models as if the treatment had occurred at that alternative time. If our results capture a genuine causal effect, we expect a pronounced treatment effect at the actual intervention time ($t=0$), and near-zero or statistically insignificant estimates at all placebo points before and after.
Figure 6 visualizes these placebo tests by plotting the precision-weighted average treatment effects and their confidence intervals across all experiments measuring animosity (see Table 5). At the actual treatment point $t=0$, which compares message content in the 30 minutes before and after the first appearance of the trigger word ([30,0[[-30,0[[ - 30 , 0 [ vs. $[0,30]$), we observe a clear and substantial effect (1.99). We then shift the treatment window rightward in 10-minute steps—first comparing [20,10[[-20,10[[ - 20 , 10 [ to $[10,40]$, and so on—until the actual intervention fully exits the analysis window at $t=30$. This generates a steady decline in estimated effects, which converge to zero, as expected if no real treatment occurred at these placebo times. Similarly, shifting the window to earlier times (e.g., comparing [60,30[[-60,-30[[ - 60 , - 30 [ to $[-30,0]$) reveals a decreasing trend in effect size, though some residual difference remains. In conclusion, given the clear peak of the effect at the timing $t=0$, we are confident in the robustness of our approach to defining the moment of the actual treatment moment.
Figure 6: Placebo treatments every 10 minutes, ranging from 30 minutes before to 30 minutes after the actual treatment point. Precision weighted average treatment effects and 95% confidence intervals across all experiments.
Figure 7: Testing the Parallel Trends Assumption with event studies. Time-wise mean of all individual event studies in purple.
5 Conclusion

### v4

@@ 3.3.2. Statistical analysis
We estimate the trigger words’ treatment effect in 60 separate experiments—12 experiments for H1 (Table 3) and 48 experiments for H2 (Table 4). For each experiment, we prepare the data in the same way. The treatment is the first mention of a trigger word in a discussion thread, i.e., there is only one well-defined treatment per thread. To test H1, we count the number of comments for each discussion thread 30 minutes before and 30 minutes after the treatment in the treatment group and the placebo treatment in the control group. We then calculate the share of comments before and after the intervention. For example, if there are 8 comments during the 30 minutes before the trigger word and 12 comments during the 30 minutes after the trigger word, we normalise the data of this discussion thread to 40% before intervention and 60% after intervention. For experiments with H2, we count the number of comments expressing animosity in a thread and then normalise the data as a share of animosity per thread before and after the trigger word.
@@
We identify the causal DiD-estimator with a two-way fixed effects model in an OLS-regression framework (Angrist and Pischke, 2009): $\text{proportion}\sim\text{trigger}+\text{treatment}+\text{interaction}$ where proportion is the per-thread-normalised target feature, and the interaction term is given by $\text{interaction}=\text{trigger}\times\text{treatment}.$ Since the DiD estimator subtracts the difference before and after intervention in the control group from the equivalent in the treatment group, no further control variables are necessary —if the parallel trends assumption holds (Roth et al., 2023). In the absence of treatment, the difference in outcomes between the treated and control groups would follow the same trajectory over time. [^5] See Section 5 for further explanations and Appendix C for full detailed analyses.
3.4. Features
We rely on comment counts to test an increase in user engagement (H1). To measure animosity (H2), we identify relevant features using various language models based on the RoBERTa architecture (Liu et al., 2019) that are specialised for social media. In Appendix A, we validate the proposed approach through a manual evaluation on a random Reddit sample across a variety of open-source transformer-based language models.
3.4.1. User engagement (Hypothesis 1)
@@
Our first hypothesis posits that trigger words increase user engagement. We rely on the number of comments to investigate this and expect an uptick in message frequency after the trigger words are mentioned for the first time in a thread.
3.4.2. Animosity (Hypothesis 2)
@@
Hypothesis 2 suggests that trigger words cause more animosity in a debate. We operationalise this with controversial comments, negative sentiment, anger, and different forms of hate speech.
@@
Controversiality. Reddit itself identifies a comment as controversial if it receives similar amounts of upvotes and downvotes. We expect trigger words to cause more controversy.
@@
Figure 2. Example: Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
@@
Sentiment analysis. We use twitter-roberta-base-sentiment-latest (Loureiro et al., 2022) to identify a comment as negative, neutral, or positive. This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal et al., 2019). Our approach builds on existing work that uses sentiment analysis to study online debates on controversial topics such as vaccination (Melton et al., 2021), politics (Guimaraes et al., 2019), and gender bias (Marjanovic et al., 2022).
@@
Negative emotions. Using twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al., 2022), we assign one or more emotions to each tweet. This model is trained on the Semeval 2018 Affect in Tweets task (Mohammad et al., 2018) and covers 11 different emotions. In our analysis, we expect to observe increased anger following a trigger word.
@@
Hate speech. Finally, we use the twitter-roberta-base-hate-multi-class model (Antypas and Camacho-Collados, 2023) to detect hate speech. It is trained on a combination of 13 hate-speech Twitter datasets and predicts hate speech directed at seven target groups. Hate speech detection is motivated by previous research revealing hateful trends in different topics (Farrell et al., 2019; Rieger et al., 2021; Walter et al., 2022). For NHS, Brexit and Vaccine, we consider any hateful comments. We focus on sexism for Feminism/Feminist, and on racism for Rwanda.
4. Results
4.1. User engagement (Hypothesis 1)
@@
According to our first hypothesis, trigger words cause more comments. The DiD estimate reveals significant differences in comment frequency between the treatment and control groups for all our trigger words (Table 3).
@@
In the treatment-only subset, the difference in user engagement before and after the trigger word is evident across all terms. It is the largest, with Rwanda, where we observe a 10%-point increase (Table 3). To identify the causal effect of the trigger word, we compare the difference in the treatment group to the difference in the control group (DiD estimate). Across all settings, we observe a significant difference in the number of comments between the treatment and control subsets, with the treatment group showing a larger increase. The lowest value is a 3.2 percentage-point increase in the contextual comparison for Brexit. The largest increase is for NHS in the contextual control group, with a 13.4% point increase.
@@
When examining comment frequency across various control groups, the impact of specific trigger words becomes more apparent. For instance, Figure 2 illustrates the distribution of comments related to the trigger Rwanda in the space-related experiment with the overall number of messages across all threads. This trigger word leads to varying levels of engagement across different Reddit communities, with a notable increase in comments in identified communities where Rwanda is more controversial. Similar patterns are also observed in the contextual and temporal control groups for Rwanda (Appendix B, Figure 4). Overall, our analysis of the influence of trigger words across different control groups indicates a clear increase in engagement for each trigger word analysed, confirming our first hypothesis.
@@
Table 4. Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
4.2. Animosity (Hypothesis 2)
@@
Our second hypothesis suggests that trigger words may increase the frequency of polarised messages and lead to more animosity in the discussion. [^6] Examples of discussions are presented in Appendix D.
@@
Figure 3. Average controversiality per thread for the treatment and control (space) groups of Rwanda.
@@
Controversiality. We analyse the average controversiality of each thread and compare the proportion of controversial comments before and after the introduction of the trigger words. Our results (Table 4) indicate that, in general, there are more controversial comments after a trigger word. The largest difference is the contextual comparison for Rwanda—a causal effect of 6.6% points. However, not all differences are statistically significant. Notably, there is no increase in controversial comments for Feminism, and in the temporal and contextual settings for Vaccine and NHS, respectively.
@@
As an example, Figure 3 shows the percentages of controversial threads related to Rwanda for the space experiment, before and after its occurrence. The data indicates a noticeable increase in thread controversiality caused by the trigger word (DiD $2.3\%$ points).
@@
Sentiment & emotion. Based on our hypothesis, trigger words tend to evoke more negative emotions in users. In our analysis, we focus on negative sentiment, in particular anger.
@@
Table 4 indicates a clear trend for the negatively charged comments of the treatment and control groups for each trigger word. We observe a significant increase in comments conveying a negative sentiment after the trigger word appears in 11 of the 12 settings tested, with the sole exception being the contextual comparison for NHS. The largest effect in sentiment (DiD $5.5\%$ points) appears for the Rwanda trigger word in the space and contextual experiments.
@@
Trigger words also cause more anger. We observe positive causal effects across all conditions, with only 3 non-significant effects. Substantively, we measure a causal effect that ranges from 1.4% points more anger (Feminist) to 5.5% points (Vaccine contextual comparison).
@@
Hate speech. As a final indicator of animosity, we also consider hate speech and abusive language, more specifically any kind of toxicity for Brexit, NHS, and Vaccine, and the more domain-specific racism-related hate speech for Rwanda and sexism-related hate speech for Feminism. As seen from Table 4, the impact of trigger words regarding the occurrence of hateful or offensive comments is not as prevalent as anger and negative sentiment. We observe significant effects in areas where we can rely on a domain-specific hate speech measure, i.e., in Feminism (space and contextual comparison) and Rwanda (space group).
5. Robustness of Results
We assess the robustness of our Difference-in-Differences (DiD) estimates through parallel trends tests and placebo treatments. The parallel trends assumption requires that, absent treatment, the treatment and control groups would follow similar trends over time. To assess this, we implement event studies that trace changes in the outcome variable across six 10-minute intervals before and after the first mention of a trigger word. As shown in Appendix C, Figure 6, pre-treatment trends remain flat and centred around zero, supporting the validity of the DiD design.
To further validate our identification strategy, we conduct placebo analyses by shifting the assumed treatment point in 10-minute increments before and after the actual intervention. For each placebo word, we re-estimate the DiD models. If our findings were artefacts of general conversational trends, we would expect similar effects with these alternative words. Instead, Figure 5 in Appendix C shows that only the actual treatment time yields a pronounced effect, while estimates at placebo times converge toward zero. Together, our checks strengthen confidence in our causal claims.
6. Conclusion